At 8:00 am the smiths left a campground, driving at 48 mi/h. At 8:20 am the Garcias left the same campground and followed the same route driving at 60 mi/h. At what time did they overtake the Smiths? this is whatr I have so far
Smiths= 48*t
Garcias= 60(1/3+t)
I 'm not even sure if thats right. Please help!!!!

Answer by sudhanshu_kmr(1152)


distance traveled by smiths in 20 minutes = 48*1/3 = 16 mi
let after x hour Garcias will overtake then...
60*x = 16 + 48*x
=> 60x - 48x = 16
=> 12x = 16
=> x = 16/12 = 4/3 hours
x = 1 hour 20 minutes...
after 1 hour 20 minutes from 08:20 am

at 9: 40 am